OpenStudy (anonymous):

A square tile has an area of 144 square inches. What is the length of one side of the tile?

jimthompson5910 (jim_thompson5910):

Area = (side length)^2 A = x^2 144 = x^2 x^2 = 144 x = sqrt(144) x = 12 So the length of one side is 12 inches.